1. Calculate the hydrogen ion concentration in moles per liter for each of these solutions: (a) a solution whose pH is 5.20, and (b) a solution whose hydroxide concentration is 3.7 x 10-9 M.
1
Expert's answer
2020-05-13T14:21:12-0400

Solution.

1.

"pH = -lg[H^+]"

"[H^+] = 10^{-pH}"

"[H^+] = 6.31 \times 10^{-6} \ M"

2.

"pOH = -lg[OH^-]"

pOH = 8.43

"pH + pOH = 14"

"pH = 14 - pOH"

pH = 5.57

"[H^+] = 2.69 \times 10^{-6} \ M"

Answer:

1.

"[H^+] = 6.31 \times 10^{-6} \ M"

2.

"[H^+] = 2.69 \times 10^{-6} \ M"